# Disentangling Representations in RNNs through Multi-task Learning

### v1

@@ B.4 Discussion
Assumptions and Limitations:
One important consideration is the assumptions made about the noise distribution and the posterior probabilities over decision boundaries. While our main derivations assume Gaussian noise for analytical tractability, we note that the key property enabling our results is that the posterior probabilities correspond to distances from the decision boundaries in a well-behaved manner (Lemma B.3). Intuitively, this requires that the posterior distribution has a single center of mass, and integrating probability over half-spaces yields a monotonic distance function. Gaussians naturally satisfy this property as maximum entropy distributions for a given variance. Extending our results to other noise distributions satisfying similar properties is an interesting direction for future work.
@@
Our theoretical results have important implications for the manifold hypothesis, which posits that real-world high-dimensional data tend to lie on or near low-dimensional manifolds embedded in the high-dimensional space [23, 15]. The key insight is that our proofs show an optimal multi-task classifier must encode an estimate of the abstract coordinates of the true underlying environment state in its latent representation. Consider the abstract space in which $\mathbf{x}^{*}$ resides, denoted $\mathcal{X}^{*}$. The injective observation map $f:\mathcal{X}^{*}\to\mathcal{X}$, where decision boundaries $y_{i}:\mathcal{X}^{*}\to\{0,1\}$ are linear. Our results imply that an optimal classifier’s latent state $\mathbf{Z}(t)$ must encode abstract cordinates in $\mathcal{X}^{*}$ rather than ambient coordinates $\mathcal{X}$.
The injective observation map $f$ aligns closely to the typical conception of a data manifold (e.g., if $f\in C^{1}$ or $f\in C^{n}$, as described in [71]). The abstract space $\mathcal{X}^{*}$ can be seen as the intrinsic coordinate system of the manifold, while $f$ maps these coordinates to the high-dimensional observation space $\mathcal{X}$. Our findings suggest that an optimal classifier will implicitly learn to invert this mapping to recover the abstract coordinates. Moreover, for natural data where the manifold hypothesis holds, the learned latent representation would plausibly capture the manifold structure, as this is essential for disambiguating noisy observations and estimating the true underlying state. The low-dimensional manifold structure is a key prior that an optimal classifier can (and in our case must) exploit to improve its performance.
@@
Figure S4: (A) $\mathbf{x}^{*}$ is noised before being transformed by injective observation map $f$, resulting in observations $f(\mathbf{X}(t))$ lying on the image of $f$ (here $f$ is a 2D folded surface).
(b) $\mathbf{x}^{*}$ is first transformed by injective observation map $f$ and noise is added afterward, resulting in observations $f(\mathbf{x}^{*})+\sigma\mathcal{N}(0,I)$ that do not lie on the image of $f$.
The curvature of the observation map $f$ and the level of noise $\sigma$ are fundamental factors influencing the extent of the degeneracy introduced by the noise after the observation map. High curvature in $f$ can make the intrinsic geometry of the data more challenging to identify (e.g., more tightly crumpled paper). Large noise levels can push observations further from the underlying manifold, similarly worsening the potential degeneracy in the observations. Interestingly, this degeneracy is a key concern in work on visual predictive coding for map learning by Gornet and Thomson [26]. They demonstrate that a predictive coding network can learn to map a Minecraft environment with visually degenerate states by integrating information to perform predictive coding along trajectories within the environment.
@@
The theoretical analysis presented in this appendix sheds light on the factors driving the emergence of disentangled representations in neural networks and their alignment with the intrinsic structure of the data. By formalizing the conditions under which learned representations recover the true underlying data manifold, our work provides a foundation for understanding the remarkable success of representation learning across diverse domains. Avenues for future research include extending our results to more general noise distributions, exploring the sample complexity of learning under different observation maps and noise levels, and empirically validating the convergence of representations across models and modalities in the context of task diversity and energy constraints.
Appendix C Supplementary figures
Table S1: Hyperparameter values for RNN training. These values apply to all simulations, unless otherwise stated. $\tau=100\textit{ms}$ was chosen as a conservative estimate of membrane time constant. $\sigma$ was varied in simulations (e.g. Figure 3b). We found that free RT tasks benefited from a higher learning rate. Other hyperparameters worked out of the box.
@@ Appendix C Supplementary figures
Figure S5: Details of learned representations and of learning. (a) Representation after the decoder. 3 out of 40 features were chosen randomly to be displayed. Compared to Figure 1d the representations wrap around non-linearly, and the quadrants are overlapping. The RNN needs to invert the non-linear mapping and remove the noise to arrive at disentangled representations. (b) Individual trial examples from network in Figure 1c. Plotting conventions same as in Figure 1d, except here every trial has been mapped to a separate color (lines and final dot). The ground truth $\mathbf{x}^{*}$ for these example trials is shown in the bottom right, and the attractors have been made transparent for better visibility. Note that these trials include noise, like the ones the network has been trained on. As can be seen, the network maintains a sense of metric distances in the 2D space: examples close in state space are also close in representation space. (c) Representation early in learning, for a network trained with 1/4 of the examples compared to Figure 1d. The representation is not abstract yet, however it is visible how the quadrants start separating and the attractors start spreading in the 2D manifold. (d) Same as in b, but for network with a delay period of 500 ms (5 darker dots at the end of trajectories). Activity remains localized after the removal of the evidence streams, maintaining a short-term memory of the joint evidence with only minor leaks. (e) Representation learned in a network trained without input noise ($\sigma=0$). Trajectories separate from the beginning, and there is no pressure to learn a 2D continuous attractor anymore. (f) Train and test errors for linear decoder for the OOD generalization task. Transparent lines correspond to different quadrants while opaque lines to the average across quadrants for one network.
Figure S6: Disentanglement and factor correlations. (a) We introduce strong correlations in the latent factors, rendering parts of the state space virtually invisible to the network during pre-training (trained for a total of 24 classification tasks). (b) Despite that, generalization performance is excellent for correlations very close to 1. Once the factors are perfectly, performance drops significantly. This implies that the network can learn a disentangled representation from correlated inputs, as long as there is some signal about the factors independently. This finding goes beyond [38] to show that the multi-task learning setting allows OOD generalization when the distribution during training the RNN itself is vastly different that the one during testing.
Figure S7: RNN activations and relation to latent variables. (a) Steady-state firing rates of hidden layer of RNN in Figure 1c, while systematically varying $x_{1}$ and $x_{2}$ from -0.5 to 0.5. Firing rates for a total of 64 neurons are plotted in 8*8 grids, one for each value of $x_{1}$ and $x_{2}$. (b) Correlation coefficient of steady-state firing rates with $x_{1}$ and $x_{2}$, respectively.
Figure S8: Representations for a multiplicative task and for continual learning. (a) Classification lines for the multiplicative task. There is a total of 48 classification lines, 12 per quadrant. (b) Top 3 PCs of RNN activity over trials. The network still learns a two-dimensional continuous attractor, however the trajectories of different quadrants separate. (c) The network learns an abstract representation when trained for the linear and multiplicative tasks in interleaved batches, and unlike in b the trajectories of different quadrants do not separate.
Figure S9: RNN simulations confirm and extend theoretical predictions. (a) OOD $r^{2}$ for free RT network required to report its estimate of $\mathbf{x}^{*}$ at different times. Maximum network $r^{2}$ matches optimal multi-task classifier theory predictions (Equation 7). (b) We now let task demands cover only half the state space, i.e. let the 24 classification lines only cover quadrants 1 and 3. (c) Top 3 PCs of RNN activity over trials for network trained in classification lines that cover half of the state space. The network still learns a two-dimensional continuous attractor, however the representation now is not disentangled, compared to Figure 1d. For example notice how the representation of the blue and green quadrants has collapsed in the absence of task demands, and how the yellow quadrant lies on a different 2D plane than the red.

### v2

@@ B.6 Discussion
Our theoretical results have important implications for the manifold hypothesis, which posits that real-world high-dimensional data tend to lie on or near low-dimensional manifolds embedded in the high-dimensional space [24, 15]. The key insight is that our proofs show an optimal multi-task classifier must encode an estimate of the disentangled coordinates of the true underlying environment state in its latent representation. Consider the disentangled space in which $\mathbf{x}^{*}$ resides, denoted $\mathcal{X}^{*}$. The injective observation map $f:\mathcal{X}^{*}\to\mathcal{X}$, where decision boundaries $y_{i}:\mathcal{X}^{*}\to\{0,1\}$ are linear. Our results imply that an optimal classifier’s latent state $\mathbf{Z}(t)$ must encode disentangled coordinates in $\mathcal{X}^{*}$ rather than ambient coordinates $\mathcal{X}$.
The injective observation map $f$ aligns closely to the typical conception of a data manifold (e.g., if $f\in C^{1}$ or $f\in C^{n}$, as described in [72]). The disentangled space $\mathcal{X}^{*}$ can be seen as the intrinsic coordinate system of the manifold, while $f$ maps these coordinates to the high-dimensional observation space $\mathcal{X}$. Our findings suggest that an optimal classifier will implicitly learn to invert this mapping to recover the disentangled coordinates. Moreover, for natural data where the manifold hypothesis holds, the learned latent representation would plausibly capture the manifold structure, as this is essential for disambiguating noisy observations and estimating the true underlying state. The low-dimensional manifold structure is a key prior that an optimal classifier can (and in our case must) exploit to improve its performance.
Relation to Autoregressive Language & Multi-Modal Transformers:
@@
Figure S8: (A) $\mathbf{x}^{*}$ is noised before being transformed by injective observation map $f$, resulting in observations $f(\mathbf{X}(t))$ lying on the image of $f$ (here $f$ is a 2D folded surface). (B) $\mathbf{x}^{*}$ is first transformed by injective observation map $f$ and noise is added afterward, resulting in observations $f(\mathbf{x}^{*})+\sigma\mathcal{N}(0,I)$ that do not lie on the image of $f$.
The curvature of the observation map $f$ and the level of noise $\sigma$ are fundamental factors influencing the extent of the degeneracy introduced by the noise after the observation map. High curvature in $f$ can make the intrinsic geometry of the data more challenging to identify (e.g., more tightly crumpled paper). Large noise levels can push observations further from the underlying manifold, similarly worsening the potential degeneracy in the observations. The reach of the manifold $f$ ([1]) can be used as an immediate loose bound for post-observation map noise $\epsilon_{2}(t)$ to ensure that the derived theorems still hold. Interestingly, this observation map degeneracy is a key concern in work on visual predictive coding for map learning by Gornet and Thomson [27]. They demonstrate that a predictive coding network can learn to map a Minecraft environment with visually degenerate states by integrating information to perform predictive coding along trajectories within the environment.
@@
The theoretical analysis presented in this appendix sheds light on the factors driving the emergence of disentangled representations in neural networks and their alignment with the intrinsic structure of the data. By formalizing the conditions under which learned representations recover the true underlying data manifold, our work provides a foundation for understanding the remarkable success of representation learning across diverse domains. Avenues for future research include exploring the sample complexity of learning under different observation maps and noise levels, and empirically validating the convergence of representations across models and modalities in the context of task diversity and energy constraints.
Appendix C Supplementary figures & tables
Table S1: Hyperparameter values for RNN training. These values apply to all simulations, unless otherwise stated. $\tau=100\textit{ms}$ was chosen as a conservative estimate of membrane time constant. $\sigma$ was varied in some simulations (e.g. Figure 5c). We found that free RT tasks benefited from a higher learning rate. Other hyperparameters worked out of the box.
@@ Appendix C Supplementary figures & tables
Figure S9: Details of learned representations and of learning. (a) Representation after the decoder. 3 out of 40 features were chosen randomly to be displayed. Compared to Figure 3d the representations wrap around non-linearly, and the quadrants are overlapping. The RNN needs to invert the non-linear mapping and remove the noise to arrive at disentangled representations. (b) Individual trial examples from network in Figure 2b. Plotting conventions same as in Figure 3d, except here every trial has been mapped to a separate color (lines and final dot). The ground truth $\mathbf{x}^{*}$ for these example trials is shown in the bottom right, and the attractors have been made transparent for better visibility. Note that these trials include noise, like the ones the network has been trained on. As can be seen, the network maintains a sense of metric distances in the 2D space: examples close in state space are also close in representation space. (c) Representation early in learning, for a network trained with 1/4 of the examples compared to Figure 3d. The representation is not disentangled yet, however it is visible how the quadrants start separating and the attractors start spreading in the 2D manifold. (d) Same as in b, but for network with a delay period of 500 ms (5 darker dots at the end of trajectories). Activity remains localized after the removal of the evidence streams, maintaining a short-term memory of the joint evidence with only minor leaks. (e) Representation learned in a network trained without input noise ($\sigma=0$). Trajectories separate from the beginning, and there is no pressure to learn a 2D continuous attractor anymore. (f) Train and test errors for linear decoder for the OOD generalization task. Transparent lines correspond to different quadrants while opaque lines to the average across quadrants for one network.
Figure S10: Angles between latent factor decoders in higher dimensions. Angles converge to 90 degrees as $N_{\text{task}}\gg D$ for RNNs, and as early as $N_{\text{task}}\geq D$ for transformers (see Section A.3 for angle estimation details). This confirms that multi-task learning leads to orthogonal, disentangled representations, in some cases even earlier than our theoretical guarantees.